Louisiana State University Health Sciences Center Administration & Finance
 

From: Ronnie Smith, Vice Chancellor, Administration and Finance
Subject: Hurricane Katrina Recovery – Update #48
Posted: August 3, 2007

The formal recommendations from FEMA on the proposal to make permanent repairs to the Lions/LSU Clinic Building have been completed and will begin to move through the various approval queues. This proposal will elevate the Cohn Learning Center to the 6th floor and elevate critical mechanical, electrical and plumbing equipment to the 2nd floor to insure that damage would be minimal should we experience another flood event at some time in the future. The proposal also includes wet and dry flood proofing measures on the 1st floor. We are working closely with FEMA to help expedite the review processes in any way possible.

We have previously forwarded a joint conceptual Section 406 Hazard Mitigation proposal for the CSRB & MEB. Individual concepts for the Allied Health/Nursing Building, the Residence Hall and the Resource Center have also been sent. The FEMA 406 Hazard Mitigation team has begun their review of the design concepts for the Resource Center, the Residence Hall and the Allied Health/Nursing Buildings. If the concepts are deemed to be acceptable and the initial cost estimates appear to indicate that the proposals are cost effective, the architects and engineers will develop formal design and development proposals with detailed costs. These will then be submitted and will form the basis for the formal review and final approval by FEMA.

The final environmental remediation and duct cleaning at the Dental School is substantially complete. There are only a small number of air handling units that have not been cleared to be turned back on. The Central Sterilization project is proceeding on schedule. The replacement of the ceiling grid and tile in the upper floors of the Clinic and Administration Buildings is substantially complete. The contractor is now working on the punch list items. Work on the repairs to the exterior parking lot lighting and parking lot controls is on schedule.

A change order to complete some additional temporary repair work on the 1st floor of the Clinic Building has recently been prepared and approved. Work is underway to construct some temporary work space for the Facility Services staff, the Supply Chain Management staff, and the Dental School’s Central Supply function. The architects are projecting to have their initial design concept for the 406 Hazard Mitigation and permanent repairs ready for LSUHSC and Facility Planning to review early this week.

Temporary repairs continue to be made at 1542 Tulane Ave. The final stages of work to restore power to the remaining portions of the building are nearing completion. Building outages occurred at the end of this week to facilitate this work. The temporary emergency generator is on site and a change order has been proposed for the construction of a foundation on which it can be mounted. The air compressor system that is required by the building system controls is being replaced. One elevator has been repaired and can be certified for normal use now that the emergency telephone lines are connected. Repairs are moving quickly on a 2nd elevator and the repair work is getting underway on two more elevators.

Improvements to the primary Student Parking Lots on the downtown campus are moving forward. Both Student Lot #2 and #3 have been paved with asphalt; replacing the old gravel surfaces. Lot #3 has been striped and the concrete bumpers have been installed. New sidewalks have been poured along the side of Lot #2, and in front of both lots. A seating area has been added to Lot #2 and will be landscaped in the future. Both lots have been reopened, although short term closures are possible to complete some remaining work. Any closures will be announced in advance.

Here are highlights of work that has been accomplished since our last update was posted two weeks ago.

Project to Make Temporary Repairs to the Dental School:

  • The final environmental remediation and duct cleaning contract is substantially complete.
    • Only a few air handling units have not yet passed the viable sample tests.
  • The project to upgrade the network infrastructure continues to move towards completion.
    • The testing of all the new network cable installed in the Clinic Building has been completed.
    • All of the new cable has been installed in the Administration Building and the testing of the cable is nearing completion.
    • The construction of the new network/telecom closets is progressing. The installation of doors and ceiling lights in the closets is well underway throughout both buildings.
    • There is still some electrical, lighting and HVAC work to be completed in the new network/telecom closets.
  • The contract to replace the ceiling tiles and grid in the Clinic and Administration Building is substantially complete.
    • The architects have prepared a punch list for this project and provided it to the contractor.
    • Work is underway to address the punch list items.
  • The contract to restore the Central Sterilization function in the 3rd floor location in the Clinic Building space is in place and the work is well underway.
    • The wall framing is complete.
    • The plumbing and electrical installation in the new Central Sterilization area is complete.
      • All that remains is to make the final connections after the sterilizers and related equipment is moved into their final positions.
    • The door frames and sheetrock have been installed in the new Central Sterilization space.
  • The repairs to the exterior site lighting and the electrical work necessary to power the parking lot controls are continuing.
    • New light poles are expected to be delivered and installed shortly.
  • Most of the door hardware for the replacement doors and frames needed for the 1st floors has been delivered and the installation work is nearing completion.
    • Panic devices are being installed on the fire exit doors.
    • When all of the hardware has been installed, the project will be complete.
  • Carpet has been installed on the 2nd and 3rd floors of the Administration Building.
    • Carpet installation in the Library space is well underway.
  • A change order proposal to replace a number of HVAC inline dampers in the buildings has been received and approved.
    • LSUHSC had a limited number of dampers on site and work has begun to install those devices.
    • The rest of the dampers are on order.
  • A change order proposal to make repairs to the compressed air line running through the utility tunnel from the Power Plant to the Clinic Building has been received and approved.
    • Work is underway on these repairs.
  • Representatives of the roof testing company completed their analysis and provided the results to the architects.
    • The test results and the architects’ recommendation to replace the roofs on the Clinic and Administration Building have been forwarded to FEMA for their consideration.
    • The architects are nearing completion of the construction and bid documents for this work.
    • If FEMA concurs with the recommendation to replace the roofs, bids for the project can be advertised.
  • Schindler Elevator has been working on completing repairs to the three passenger elevators in the Administration Building.
  • Both of the sections of the new cooling tower were lifted onto the roof of the Power Plant and set on the prepared steel structure.
    • The final connections were made to the new cooling tower and the change over to the new unit was successfully completed.
    • The demolition of the old cooling tower has been completed.
  • The removal of the flood damaged fire brick from the incinerator is nearly complete.
    • Refractory bricks were delivered for the incinerator and the installation of the new bricks will commence following the removal of the existing bricks.
  • Work on the Change Order to create temporary work spaces for Supply Chain Management, Facility Services and Central Supply on the 1st floor of the Clinic Building has begun.
    • Electricians were on site and started work.
    • Demolition of old duct work, electrical wiring, etc. is underway.
    • Duct removal is temporarily on hold until the asbestos wrap used to seal portions of the old duct work is abated.
    • Temporary lighting is being installed.
    • The air handling units have been disconnected and moved to their temporary locations.
    • The rough in for the new lights, electrical outlets, etc. is continuing.
  • The demolition of the AV island in Auditorium C was completed with the removal of the fan coil unit and the repositioning of the electrical service.

Project to Make Temporary Repairs to 1542 Tulane Ave. :

  • Bell Roofing successfully removed the shed and patched the foundation on the 9th floor roof.
    • They installed a 2nd layer of roofing over the concrete foundation on the roof.
  • Ordes is pulling new wiring and is working on the junction box connections between the new high-voltage switchgear in the basement and the 4th floor switchgear.
    • They completed the junction box connections between the new high-voltage switchgear and the 4th floor riser.
    • The 4th floor riser has been energized the 4th floor electrical riser and the contractor verified that the power was being fed to all the areas that it supplies.
      • A missing fuse along with spares for the 4th floor riser switch gear are on order. This will provide power for the upper floors on the Tulane side of the Bldg.
    • Ordes completed the swap out of the 480V transformer with the proper 280V transformer and has re-energized the transformer.
      • This transformer provides power for the panels PLL6 and 7 which powers the lights in the old section of the building near the Entergy Garage.
      • Ordes Electrical finished closing up the PL4 panel in the vault.
    • They also started running conduit to the L & M Bldg Parking controls.
  • Stratos Elevator moved elevator equipment to the penthouse of Elevators #7 and #8.
    • Stratos received the door operators, door hardware, and the door tracks for Elevators #7 and # 8 (the Tulane side elevators).
    • They have removed the flood damaged cab walls of Elevator #2.
      • The flooring of Elevator #2 has been removed and new sub-flooring has been installed.
      • They are continuing to work on rebuilding the Elevator #2 cab.
      • Stratos began removing the old safeties for the #2 elevator. All of the cab work is complete except the canopy.
    • They began installing the controllers for the #7 and #8 elevators.
  • Siemens is on site removing the damaged fire alarm wiring and conduit it the basement.
    • Work is underway on the 2nd, 3rd, 4th, and 5th floors.
  • Ordes replaced the brick pavers across the drive after running the conduit over to the L&M lot and began cutting and breaking out the concrete to run the conduit over to the parking booth.
    • Conduit has been run underneath the drive over to the parking booth in the L&M lot at the Medical School Building. The pavers have been replaced and the concrete is being replaced.

Downtown Campus Landscaping Project:

  • Irrigation pipe was installed on the Bolivar St. side of the Allied Health/Nursing Building under the overhang where the gravel was removed.
    • The irrigation zone on the Bolivar St. side of the Allied Health/Nursing Building under the overhang is fully operational, but is currently using a temporary water source. 
  • Crews completed the installation of the irrigation zone on the Bolivar St. side of Allied Health/Nursing under the building overhang.
    • They adding new topsoil to the same area and grading it in preparation for planting. 
    • Crews completed planting all of the variegated Algerian ivy on the Bolivar St. side of the Allied Health/Nursing Building (under the overhang and the walkway) that was on hand. Additional plants are on order for this area.
  • Crews continued trenching for the installation of the irrigation main and related wiring on the S. Roman St. side of Allied Health/Nursing and the MEB and on the Perdido St. side of the MEB.
    • Electrical wire was installed for five irrigation valves on the Bolivar St. side of the building and for two of the valves on the S. Roman St. side of the Allied Health/Nursing Building and the MEB.
  • All of the short sleeves on the Perdido St. side of the MEB have been installed.
  • Crews finished trenching for the irrigation main and wires on the S. Roman St. and Perdido St. sides of the MEB.
  • The crew connected the new irrigation main for the Allied Health/Nursing Building and the MEB to a backflow preventer on the S. Roman St. side of the buildings. 
    • They also installed the main piping and three valves in same area. 

MEB Cafeteria Project:

  • The temporary wall separating the construction area from the hallway has been removed.
  • The drywall has been taped and floated.
  • The installation of drywall on the ceiling grid is well underway.
  • The sprinkler system has been installed and successfully passed a pressure test.
  • The first of the new lighting circuits were energized.

Student Parking Lot Re-Surfacing:

  • Student Lot #2:
    • The contractor striped Student Lot #2 and installed the new concrete wheel stops.
    • Lot #2 was re-opened for use.
    • Some work remains to be completed which will likely require closing the lot or portions of the lot for short periods of time.
      • Notices will be sent out in advance of any addition closings.
      • Work that remains includes the installation of some new light poles and a curb to control the floor of traffic through the entrance and exit.
  • Student Lot #3:
    • The contractor reinstalled the island gate controllers in Student Lot #3.
    • The contractor also set the posts for the card readers and has installed the gate arms & card readers.
    • Work in Lot #3 is substantially complete.
      • Work remaining includes the installation of a barrier or curb to control the flow of traffic through the entrance and exit, and the replacement of a light pole.
    • Lot #3 has been re-opened for use.

MEB Roof Replacement:

  • Bids were opened on 7/24/07 as scheduled.
  • The bids have been evaluated and the low, responsive bidder has been identified.
  • The architects and LSUHSC have forwarded their award recommendations to the Office of Facility Planning and Control who will be awarding the contract.
  • We would hope that work can begin within the next few weeks.

CSRB 1st Floor Vivarium:

  • Arc Mechanical completed the process of extending the piping in preparation for elevating the Animal Care chiller.

MEB Duct Cleaning:

  • In our continuing effort to improve the quality of our facilities, on Monday, July 30th we began a project to clean the air ducts throughout the Medical Education Building. This cleaning will remove buildup of particulate matter, preventing airborne contamination.
    • In addition to the cleaning, the ducts will be coated with an antimicrobial polymer. This coating is a non-toxic polymer that will lock down any remaining material and help inhibit potential mold growth.
    • The cleaning activities will take place at night from 6:00 pm to 6:00 am. A facilities representative is assigned to the project and will assist in coordinating the cleaning activities and grant access to the areas on each floor.
  • The tentative schedule for cleaning will be as follows:
    • 4th floor - July 30 – August 2
    • 3rd floor - August 3 – 7
    • 2nd floor - August 8 – 11
    • 5th floor - August 13 – 16
    • 6th floor - August 17 – 21
    • 7th floor - August 22 – 25

Campus Security:

  • One of the initiatives to enhance the security of the campus is to improve on the exterior lighting. A project is underway to replace all of the old lights in the Roman St. Garage with new lights that are both brighter and more energy efficient.
    • The result will be brighter light throughout the parking garage, without any increase in energy consumption.
    • Lights have been replaced on the majority of the 5th and 6th floors of the garage, as the crew began working at the top of the structure and are working their way down.

We encourage you to follow the details of our progress through the daily updates which are posted on the web site.